OpenStudy (anonymous):

Tim has 30 jellybeans, with equal numbers of three flavors—raspberry, pineapple, and orange. He likes raspberry the best. So far, Tim has eaten 6 orange and 4 pineapple jellybeans. What are the chances that the next one he picks will be raspberry?

OpenStudy (ankit042):

initial state pineapple call it p =10 raspberry call it r = 10 orange call it o = 10 after eating p = 10-4 =6 o = 10-6 =4 r = 10 so now total jelly beans are 20 p(r) = number of r/total 10/20

OpenStudy (ankit042):

I need a jellybean now lol

OpenStudy (anonymous):

thank you !!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

R = 10 - 0 = 10 left P = 10 -4 = 6 left O = 10 - 6 = 4 left 20 left, of which 10 are Raspberry, thus 10/20 probability. Oops @ankit042 has beaten me to it.

OpenStudy (ankit042):

@douglaswinslowcooper there are enough jellybeans we can share them lol

OpenStudy (anonymous):

thanks to both of you :)

OpenStudy (anonymous):

Thank you both, too. Yummy.

OpenStudy (anonymous):

verified proof: http://prntscr.com/7cu960